OpenStudy (anonymous):

Let h(x)=5−3x^3, h'(3)=

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

could you first find h'(x) by differentiating h(x) ?

OpenStudy (anonymous):

Use this to find the equation of the tangent line to the curve y=5−3x3 at the point (3−76) and write your answer in the form: b) y=mx+b, where m is the slope and b is the y-intercept.

OpenStudy (anonymous):

yes inverse function isnt

OpenStudy (anonymous):

i got y=(-x+5)?

hartnn (hartnn):

inverse ? no... can you differentiate 5- 3x^3 ?

OpenStudy (anonymous):

mmmm.. no. i have no idea. :/

OpenStudy (anonymous):

aren't you just looking for the derivative of h?

hartnn (hartnn):

if you haven't learned derivatives, how are you attempting this question ?

hartnn (hartnn):

anyways, you'll need this : \(\dfrac{d}{dx}x^n=nx^{n-1}\)

OpenStudy (anonymous):

im confused already. I want to learn.

OpenStudy (anonymous):

how do i apply this equation to the question

hartnn (hartnn):

\(\dfrac{d}{dx}x^n=nx^{n-1} \\ \dfrac{d}{dx}x^3=...?\)

OpenStudy (anonymous):

nx^3-1?

hartnn (hartnn):

note that n=3

OpenStudy (anonymous):

yes.

hartnn (hartnn):

so, \(\dfrac{d}{dx}x^3=3x^{2}\) gott his ?

OpenStudy (anonymous):

,mmmmm so answer is 3x^2? how come?

hartnn (hartnn):

i'll use a shorthand notation (x^3)' = 3x^2

hartnn (hartnn):

thats not final answer...just (x^3)'

OpenStudy (anonymous):

ok! i got it. so (x^3)'=3x^2

hartnn (hartnn):

also derivative of constant =0 so, (5)'=0

OpenStudy (anonymous):

mmm ok.

hartnn (hartnn):

so, \((5-3x^3)'=0-3(3x^2)=...?\)

OpenStudy (anonymous):

im taking note hold on

OpenStudy (anonymous):

why its zero, is it beacuse constant?

hartnn (hartnn):

yes, any constant (say c) can be written as , \(cx^0\) and if you use the formula for x^n, \((cx^0)'=c (0x^{0-1})=0\) got this ?

OpenStudy (anonymous):

YES! i think so. im taking note hold on please

hartnn (hartnn):

ok.

OpenStudy (anonymous):

Ok, so the 9x^2?

hartnn (hartnn):

-9x^2

OpenStudy (anonymous):

i forgot to write -.

OpenStudy (anonymous):

and whats the next step?

hartnn (hartnn):

so, h'(x) = -9x^2 to get h'(3) just put x=3 in that!

OpenStudy (anonymous):

ohh cool so -81?

OpenStudy (anonymous):

i got your message, i will wait here

hartnn (hartnn):

yes, -81 is correct. that will be the slope of your line, m=h'(3)=-81 now you just need b = y-intercept=....

hartnn (hartnn):

to get that, when x= 3, what is h(3) =... ?

OpenStudy (anonymous):

-3^3?

hartnn (hartnn):

where did 5 go ?

hartnn (hartnn):

note that this time its h(3) and not h'(3)

hartnn (hartnn):

so, 5-3(3^3)

OpenStudy (anonymous):

Oh why its not h'(3)?

OpenStudy (anonymous):

i got it.

hartnn (hartnn):

y= h(x) = 5-3x^3 when x=3, we need to find y thats why.

hartnn (hartnn):

that rhymed :P

OpenStudy (anonymous):

I see

OpenStudy (anonymous):

:)

hartnn (hartnn):

so, h(3)=.. ?

OpenStudy (anonymous):

h(3)=5-3(3^3)?

OpenStudy (anonymous):

76?

hartnn (hartnn):

5-3*27 must be negative..

OpenStudy (anonymous):

ops-76?

hartnn (hartnn):

oh.. is you '-' key of your keyboard not working? :P just kidding.. so, you have m= -81 , x=3,y=-76 just plug in y=mx+b and find b !

OpenStudy (anonymous):

so -167? since its -76=-243*3+b?

hartnn (hartnn):

y=mx+b -76=-81(3) +b b= 81*3-76 = 167 O.o now you have an extra -

OpenStudy (anonymous):

so no negative this time?

hartnn (hartnn):

yeah...just 167 so your equation will be y= -81x+167

OpenStudy (anonymous):

I entred it but it told me wrong :(

hartnn (hartnn):

let me go through all steps again...

OpenStudy (anonymous):

Use this to find the equation of the tangent line to the curve y=5−3x^3 at the point (3−76) and write your answer in the form:

hartnn (hartnn):

couldn't find any error :\ y= -81x+167 must have worked...

OpenStudy (anonymous):

I am so sorry i entred 160!

OpenStudy (anonymous):

it worked! thank yu!!

hartnn (hartnn):

ok. no problem. and welcome ^_^

OpenStudy (anonymous):

A. Find the average velocity for the time period beginning when t=3 and lasting